If the point(3,-2) lies on the circle defined by x^2+y^2+mx-4y-3=0, what is the value of m???
Answer by stanbon(75887)

If the point(3,-2) lies on the circle defined by x^2+y^2+mx-4y-3=0, what is the value of m???
----------------------------
Substitute for x and for y to solve for m:
3^2 + (-2)^2 +3m -4*-2 -3 = 0
9 + 4 + 3m + 8 -3 = 0
3m = -18
m = -6
===============
Cheers,
Stan H.